A truck travelling at 22.5m/s decelerates at 2.27m/s2. How much time does it take for truck to stop?


1
Expert's answer
2021-04-19T17:06:19-0400

The time it takes is


"t=\frac{v_f-v_i}{a}=\frac{0-22.5}{-2.27}=9.911\text{ s}."